OpenStudy (anonymous):

A car travels at constant speed for 2 days and covers 1700 miles. What is the speed in centimeters/minute? Use the conversion factors on page 55.

OpenStudy (anonymous):

Was there something wrong with my previous answer? If so, let me know and I'll try to help.

OpenStudy (yttrium):

From \[v = \frac{ s }{ t }\]] \[v = \frac{ 1700mi }{ 2 days } (\frac{ 1day }{ 24hr })(\frac{ 1hr }{ 60\min })(\frac{ 1.603km }{ 1mi })(\frac{ 100000cm }{ 1km })\] Then you will arrive at an answer =)